# Respective Roles of Electron-Phonon and Electron-Electron Interactions in the Transport and Quasiparticle Properties of SrVO

### v1

@@ Abstract
The spectral and transport properties of strongly correlated metals, such as SrVO (SVO), are widely attributed to electron-electron (-) interactions, with lattice vibrations (phonons) playing a secondary role. Here, using first-principles electron-phonon (-ph) and dynamical mean field theory calculations, we show that -ph interactions play an essential role in SVO: they govern the electron scattering and resistivity in a wide temperature range down to 30 K, and induce an experimentally observed kink in the spectral function. In contrast, the - interactions control quasiparticle renormalizations and low temperature transport, and enhance the -ph coupling. We clarify the origin of the near temperature dependence of the resistivity by analyzing the - and -ph limited transport regimes. Our work disentangles the electronic and lattice degrees of freedom in a prototypical correlated metal, revealing the dominant role of -ph interactions in SVO.
@@
Strontium vanadate, SrVO (SVO), is a perovskite oxide widely studied as a prototypical correlated metal [1, 2, 3]. Experiments have measured transport and spectral functions in detail in SVO, owing to advances in growth of clean samples [4, 5] and characterization by angle-resolved photoemission spectroscopy [6, 7, 8].
@@
In addition, transport measurements have found a near -dependent resistivity in broad temperature ranges below 300 K [10, 11, 12, 13].
@@
We calculate the electronic structure, phonon dispersions, and -ph coupling using DFT and density functional perturbation theory (DFPT) with the Quantum Espresso package [30, 31, 32, 33]. We use the experimental lattice parameter of 3.842 Å [2, 13] and project the electronic structure onto the t -orbitals of vanadium [34].
@@
The - self-energy is obtained with DFT+DMFT using the TRIQS code with a continuous-time quantum Monte Carlo solver [36, 37, 38, 39, 40] and Padé analytical continuation [37].
@@
As shown in Fig. 1(a), our DFT calculations predict a bandwidth of 2.5 eV for the t electronic bands, which is renormalized by a factor by DMFT, in agreement with experiments [6] and previous DMFT results [14, 16]. In the temperature range we study (115390 K), the imaginary part of the - self-energy, , follows a Fermi liquid behavior.
@@
which involve distortions of the VO octahedra, such as Jahn-Teller modes.
@@
The dominant role of - interactions on quasiparticle renormalization in SVO, despite their small effect on low energy scattering, can be rationalized using the Kramers-Kronig relations [42, 44]:
@@
at energy for electron band and momentum . Here, is the DFT band energy, is the Fermi energy, and is the electron self-energy. Following our previous work [29], in separate calculations we compute this Green’s function using the self-energy from DMFT - interactions, the lowest-order self-energy from -ph interactions, and their sum 222The -ph self-energy calculation uses the DFT electron propagator. Unlike in SrRuO [29], using the DMFT electron propagator has a negligible effect in SVO., obtaining corresponding spectral functions capturing different combinations of interactions [Fig. 2(c)-(e)].
@@
Numerous experiments have measured a near temperature dependence of the resistivity in SVO below 300 K [46, 47, 2, 4, 11, 48, 49, 50, 13].
@@
Recently, the growth of ultra-clean samples has enabled detailed measurements of the resistivity with reduced defect scattering [4, 13]. In these samples, Ahn et al. [13] find a near- resistivity below 25 K and between about 100300 K, together with a stronger than temperature dependence at intermediate temperatures.
@@
The resistivity for different combinations of interactions is shown in Fig. 3(a) and compared with experimental data [11, 13].
@@
The temperature dependence of the -ph limited resistivity is analyzed in more detail in Fig. 3(b), where we show our results on a log-log plot and compare them with experiments [11, 13]. In that plot, the resistivity is computed with the full (iterative) solution of the Boltzmann transport equation (BTE) [35] to include backscattering and improve the treatment of acoustic phonons. The computed -ph limited resistivity follows a temperature dependence between 100200 K, in excellent agreement with the dependence found in experiments in that temperature range [11, 13], and falls to at 300 K. We identify the origin of this nearly temperature trend of the -ph limited resistivity by analyzing the contribution of different phonon modes. Our calculations show that the increasing contribution of strongly coupled optical phonons at higher temperatures is responsible for the dependence of the resistivity between 100200 K (see SM [41]).
@@
We find a clear crossover between 2030 K from -ph to - dominated transport. The -ph limited resistivity becomes much smaller than the - limited resistivity below 20 K, showing that - scattering governs transport at low temperature. This result indicates that - interactions are the origin of the resistivity observed experimentally below 25 K, with the -ph contribution causing deviations from a behavior above 25 K. Note that our DMFT resistivity underestimates the experimental value below 25 K by a factor of 3. We attribute this discrepancy to limitations in the description of electronic interactions in the single site DMFT impurity model [52, 29]. A more accurate treatment may require going beyond single-site DMFT [53, 54].
@@
Strong electronic interactions are known to enhance -ph coupling in correlated metals.
@@
To study the role of correlations in SVO, we compute the -ph interactions using Hubbard-corrected DFPT (DFPT+) [33, 32], which captures the strong local interactions between -orbitals in a static approximation and accounts for the resulting change in orbital polarization response. We use a Hubbard- parameter of 3 eV, which provides orbital polarization and band structure responses to phonon perturbations similar to a calculation including accurate values of both Hubbard- and Hund’s interactions (see SM [41]).
Adding the Hubbard correction has a small effect on the phonon dispersions, but it enhances the -ph interactions, as shown in Fig. 4(a). The enhancement is mode-dependent and is generally higher for strongly-coupled optical phonons involving VO distortions.
@@
Notably, the -ph limited resistivity increases by 35% at room temperature [Fig. 4(b)], bringing the resistivity computed with both -ph and - interactions into very good agreement with experiments. For example, the computed resistivity at 290 K is 16 cm, versus an experimental value of 1925 cm [11, 13].
@@
Acknowledgements.— We thank Andrew Millis and Jennifer Coulter for helpful discussions. D.J.A. is supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2139433. This work was also supported by the National Science Foundation under Grant No. OAC-2209262, which provided for code development. D.J.A. and M.B. were partially supported by the AFOSR and Clarkson Aerospace under Grant No. FA95502110460. J.-J.Z. acknowledges support from the National Key R&D Program of China (Grant No. 2022YFA1403400). J.M. is supported by the Slovenian Research Agency (ARIS) under Grants No. P1-0044 and J1-2458. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0026831. The Flatiron Institute is a division of the Simons Foundation.

### v2

@@ Abstract
The spectral and transport properties of strongly correlated metals, such as SrVO3 (SVO), are widely attributed to electron-electron (-) interactions, with lattice vibrations (phonons) playing a secondary role. Here, using first-principles electron-phonon (-ph) and dynamical mean field theory calculations, we show that -ph interactions play an essential role in SVO: they govern the electron scattering and resistivity in a wide temperature range down to 30 K, and induce an experimentally observed kink in the spectral function. In contrast, the - interactions control quasiparticle renormalizations and low temperature transport, and enhance the -ph coupling. We clarify the origin of the near temperature dependence of the resistivity by analyzing the - and -ph limited transport regimes. Our work disentangles the electronic and lattice degrees of freedom in a prototypical correlated metal, revealing the dominant role of -ph interactions in SVO.
@@
Strontium vanadate, SrVO3 (SVO), is a perovskite oxide widely studied as a prototypical correlated metal [1, 2, 3]. Experiments have measured transport and spectral functions in detail in SVO, owing to advances in growth of clean samples [4, 5] and characterization by angle-resolved photoemission spectroscopy [6, 7, 8].
@@
In addition, transport measurements have found a near -dependent resistivity in broad temperature ranges below 300 K [10, 11, 12, 13, 14].
@@
We calculate the electronic structure, phonon dispersions, and -ph coupling using DFT and density functional perturbation theory (DFPT) with the Quantum Espresso package [31, 32, 33, 34]. We use the experimental lattice parameter of 3.842 Å [2, 13] and project the electronic structure onto the t2g -orbitals of vanadium [35].
@@
The - self-energy is obtained with DFT+DMFT using the TRIQS code with a continuous-time quantum Monte Carlo solver [37, 38, 39, 40, 41, 42] and Padé analytical continuation [38]. We use Hubbard-Kanamori interactions with eV and eV to produce spectra in agreement with experiment [6, 7, 8].
@@
As shown in Fig. 1(a), our DFT calculations predict a bandwidth of 2.5 eV for the t2g electronic bands, which is renormalized by a factor by DMFT, in agreement with experiments [6] and previous DMFT results [15, 17]. In the temperature range we study (115390 K), the imaginary part of the - self-energy, , follows a Fermi liquid behavior.
@@
which involve distortions of the VO6 octahedra, such as Jahn-Teller modes.
@@
The dominant role of - interactions on quasiparticle renormalization in SVO, despite their small effect on low energy scattering consistent with a Fermi liquid, can be rationalized using the Kramers-Kronig relations [44, 46]:
@@
at energy for electron band and momentum . Here, is the DFT band energy, is the Fermi energy, and is the electron self-energy. Following our previous work [30], in separate calculations we compute this Green’s function using the self-energy from DMFT - interactions, the lowest-order self-energy from -ph interactions, and their sum 222The -ph self-energy calculation uses the DFT electron propagator. Unlike in Sr2RuO4 [30], using the DMFT electron propagator has a negligible effect in SVO., obtaining corresponding spectral functions capturing different combinations of interactions [Fig. 2(c)-(e)].
@@
Numerous experiments have measured a near temperature dependence of the resistivity in SVO below 300 K [48, 49, 2, 4, 11, 50, 51, 52, 13, 14].
@@
Recently, the growth of ultra-clean samples has enabled detailed measurements of the resistivity with reduced defect scattering [4, 13, 14]. In these samples, Ahn et al. [13] and Brahlek et al. [14] find a near- resistivity below 25 K and between about 100300 K, together with a stronger than temperature dependence at intermediate temperatures.
@@
The resistivity for different combinations of interactions is shown in Fig. 3(a) and compared with experimental data [11, 13, 14].
@@
The temperature dependence of the -ph limited resistivity is analyzed in more detail in Fig. 3(b), where we show our results on a log-log plot and compare them with experiments [11, 13, 14].
In that plot, the resistivity is computed with the full (iterative) solution of the Boltzmann transport equation (BTE) [36] to include backscattering and improve the treatment of acoustic phonons.
The computed -ph limited resistivity follows a temperature dependence between 100200 K, in excellent agreement with the dependence found in experiments in that temperature range [11, 13, 14], and falls to at 300 K. We identify the origin of this nearly temperature trend of the -ph limited resistivity by analyzing the contribution of different phonon modes. Our calculations show that the increasing contribution of strongly coupled optical phonons at higher temperatures is responsible for the dependence of the resistivity between 100200 K (see SM [43]).
@@
We find a clear crossover between 2030 K from -ph to - dominated transport. The -ph limited resistivity becomes much smaller than the - limited resistivity below 20 K, showing that - scattering governs transport at low temperature. This result indicates that - interactions are the origin of the resistivity observed experimentally below 25 K, with the -ph contribution causing deviations from a behavior above 25 K. Note that our DMFT resistivity underestimates the experimental value below 25 K by a factor of 2-3. We attribute this discrepancy to limitations in the single site DMFT impurity model, as transport can be sensitive to the methods used to downfold orbitals and effective interactions [54, 55, 56, 30], the lack of non-local contributions to the DMFT self-energy [57], and corrections to the current-current vertex [58, 59, 60, 61].
@@
Strong electronic interactions are known to significantly modify -ph interactions [62, 63, 64, 65]. In correlated metals, -ph coupling is often enhanced.
@@
To study the role of correlations in SVO, we compute the -ph interactions using Hubbard-corrected DFPT (DFPT+) [34, 33]. DFT+ can be viewed as a static approximation of DFT+DMFT [66]; DFT+DMFT -ph coupling calculations are not currently possible. Like DFT+DMFT, DFT+ captures the strong local interactions between -orbitals and the resulting change in orbital polarization response, which modifies -ph coupling. We use a Hubbard- parameter of 3 eV in the atomic orbital basis, which is somewhat empirical but provides orbital polarization and band structure responses to phonon perturbations similar to a calculation including a Hubbard- parameter from linear response and a reasonable value of Hund’s . The orbital occupation responses are also similar to those found using DMFT with the same orbitals and interaction parameters used throughout (see SM [43])
Adding the Hubbard correction has a small effect on the phonon dispersions, but it enhances the -ph interactions, as shown in Fig. 4(a). The enhancement is mode-dependent and is generally higher for strongly-coupled optical phonons involving VO6 distortions.
@@
Notably, the -ph limited resistivity increases by 35% at room temperature [Fig. 4(b)], bringing the resistivity computed with both -ph and - interactions into very good agreement with experiments. For example, the computed resistivity at 290 K is 16 cm, versus an experimental value of 1925 cm [11, 13, 14].
@@
Acknowledgements.— We thank Andrew Millis, Jennifer Coulter, and Roman Engel-Herbert for helpful discussions. D.J.A. is supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2139433. This work was also supported by the National Science Foundation under Grant No. OAC-2209262, which provided for code development. D.J.A. and M.B. were partially supported by the AFOSR and Clarkson Aerospace under Grant No. FA95502110460. J.-J.Z. acknowledges support from the National Key R&D Program of China (Grant No. 2022YFA1403400). J.M. is supported by the Slovenian Research Agency (ARIS) under Grants No. P1-0044 and J1-2458. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award NERSC DDR-ERCAP0026831. The Flatiron Institute is a division of the Simons Foundation.